OpenStudy (anonymous):

what is 64-x/8-sqrt of x when x approaches 64?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 64} (64-x)/(8-\sqrt{x})\] You can see that upside (64-x) Is a difference of squares And you can factorize (64-x)=(8-√x)(8+√x) So i will leave us with \[\lim_{x \rightarrow 64} (8-√x)(8+√x)/(8-\sqrt{x})\] Cancel up and down. (8-√x) And evaluate the limit, and it will be. 16

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

You are welcome.